it takes 365 days for the earth to complete one revolution around the sun. Calculate period, frequency and angular frequency of earth as it revolves around the sun
1
Expert's answer
2021-10-06T08:15:49-0400

(a) We can find the period as follows:


"T=365\ days\cdot\dfrac{24\ h}{1\ day}\cdot\dfrac{3600\ s}{1\ h}=3.15\cdot10^7\ s."

(b) We can find the frequency as follows:


"f=\dfrac{1}{T}=\dfrac{1}{3.15\cdot10^7\ s}=3.2\cdot10^{-8}\ Hz."

(c) We can find the angular frequency as follows:


"\omega=2\pi f=2\pi\cdot3.2\cdot10^{-8}\ Hz=2.0\cdot10^{-7}\ \dfrac{rad}{s}."